INFRARED SPECTRA OF MOLECULAR IONS FORMED FROM PF_{3} AND PF_{5} AND TRAPPED IN A NEON MATRIX

\underline{CATHERINE L. LUGEZ} and MARILYN E. JACOX, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899; KARL K. IRIKURA, Physical and Chemical Properties Division, National Institute of Standards and Technology, Gaithersburg, MD 20899.

When PF_{3} or PF_{5} is subjected to Penning ionization and photoionization by excited neon atoms and their resonance radiation (16.6-16.8 eV) and the products are rapidly frozen in an excess of solid neon, the resulting solid deposit shows prominent infrared absorptions which can be assigned to cation and anion products, including PF_{3}^{+}, PF_{4}^{+}, and PF_{4}^{−}. Because phosphorus and fluorine each possess only one stable isotope, consideration of the spectra of related systems and of the results of ab initio calculations is important in making spectral assignments. Processes which lead to the stabilization of the observed ionic species will be discussed.